OpenStudy (bibby):

Approaching a molality problem. The density of a 1.95M solution of KBr in water is 1.16 g/mL. Find the molality of this solution. (FW KBr = 119amu) The way I'd start is see that there's 1.95mol KBr (solute)/1L of solution. I'm not sure how to approach past this point or use the density.

OpenStudy (aaronq):

molality = moles of solute/Kg of solvent find how much water you have in Kg

OpenStudy (bibby):

I don't know how to go about that. Would I multiply the density by !L and divide by 1000?

OpenStudy (anonymous):

1.95 * 119 = Mass 1.16 g/mL = 1160 g/L Mass of Solvent = 1160 - 119*1.95 After getting this Convert it into kg Then Molality = 1.95 / mass of Solvent